OpenStudy (anonymous):

Find S20 when t1 = 4 and d = 5

OpenStudy (anonymous):

Now you have to use formula

OpenStudy (anonymous):

tn=t1+(n-1)d

OpenStudy (anonymous):

\(S_{n}=\frac{n}{2} (2 t_1 + (n-1)d)\) sub n=20, and everything you've got there. do u need further help?

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

S(20=20/2(2(4)+(20-1)5)

OpenStudy (anonymous):

And how would I work this out @Shadowys

OpenStudy (anonymous):

um,simplify this then. \(\frac{20}{2}[(2)(4)+(20-1)5]=10[8+19(5)]\) what follows can be done by manual arithmetic or a calculator.

OpenStudy (anonymous):

I got 1030

OpenStudy (anonymous):

yup, that's what I got too. :)

OpenStudy (anonymous):

Cool